OpenStudy (anonymous):

A 14 foot ladder is leaning on a wall, slowly sliding down. If the base of the ladder is sliding away from the wall at 1.3 ft/second, how fast is the top sliding vertically along the wall when the bottom of the ladder is 9 feet from the wall? @Calculus1

OpenStudy (amistre64):

we need a relation that has the legs of a right triangle in it id assume

OpenStudy (amistre64):

the pythag might be useful

OpenStudy (amistre64):

wall^2 + floor^2 = 9 2w w' + 2f f' = 0 2w w' = -2f f' w' = -2f f'/2w w' = -f f'/w

OpenStudy (anonymous):

with that being done i got 16.6

OpenStudy (amistre64):

w' = -f f'/w w' = -9 (1.3)/sqrt(14^2 - 9^2) = -1.091..

OpenStudy (anonymous):

pyth. theorem c =16.6

OpenStudy (amistre64):



OpenStudy (amistre64):

i get a little over 10 for the wall height

OpenStudy (anonymous):

i got 10.7 so yea.... idk how to approach this problem .. is it like the other one ?

OpenStudy (amistre64):

they are all the same; once you have a relationship that ties it all together, its just a metter of fillin gin the pieces afterwards

OpenStudy (anonymous):

A balloon is being inflated at a rate of 10 cm3/sec. At what rate is the radius of the balloon changing when the radius is 20 cm?

OpenStudy (anonymous):

i got 1256 and its wrong

OpenStudy (amistre64):

volume and raduis need to be related

OpenStudy (amistre64):

\[V_{sphere}=\frac{4}{3}pi\ r^3\]right?

OpenStudy (anonymous):

volume = 4/3pi 3r^2

OpenStudy (anonymous):

yea thats what i qot

OpenStudy (amistre64):

\[V = \frac{4}{3}\pi\ r^3\] \[V' = 4\pi\ r^2\ r'\] \[\frac{V'}{4\pi\ r^2} = \ r'\]

OpenStudy (amistre64):

V' = 10, and r=20

OpenStudy (anonymous):

.002?

OpenStudy (amistre64):

thats a good apporx yeah

OpenStudy (anonymous):

A small spherical balloon is being inflated at the rate of 30 cm3/sec. What is the volume of the balloon 45 seconds after inflation begins? do i jus multiply 30 and 45

OpenStudy (amistre64):

i would

OpenStudy (anonymous):

i now have to find the diameter of the balloon and ik diameter id r^2